OpenStudy (anonymous):

How does the graph of the line change as you make b larger but keep m the same? A. The line beomes less steep. B. The line becomes steeper. C. The line moves upward. D. The line moves downward.

OpenStudy (solomonzelman):

For example y=3x+5 and y=3x+9 You can see that the slope is the same , but the "b" of the second line is 4 units bigger. This means we shifted the line 4 units up. The diagonal is exactly same direction in both lines, y=3x+5 and y=3x+9 Also, here are the rules for the shifts (with a different example). \(\large\color{ blue }{\large {\bbox[5pt, lightblue ,border:2px solid white ]{ \large\text{ }\\ \begin{array}{|c|c|c|c|} \hline~~~~~~~~~~~~~~~~~~~~~~~~~~~\textbf{Shifts}~~~~~~~~~~~~~~~~~~~~~~~~~~~&~\rm{c~~~units~~~~} \\ \hline \\f(x)= ∛x ~~~ ⇒ ~~~ f(x)= \sqrt[3]{x \normalsize\color{red}{ -~\rm{c}} } &~\rm{to~~the~~right~} \\ \text{ } \\ f(x)= ∛x ~~~ ⇒ ~~~ f(x)= \sqrt[3]{x \normalsize\color{red}{ +~\rm{c}} } &~\rm{to~~the~~left ~} \\ \text{ } \\ f(x)= ∛x ~~~ ⇒ ~~~ f(x)= ∛x \normalsize\color{red}{ +~\rm{c} } &~\rm{up~} \\ \text{ } \\ f(x)= ∛x ~~~ ⇒ ~~~ f(x)= ∛x \normalsize\color{red}{ -~\rm{c} } &~\rm{down~} \\ \\ \hline \end{array} }}}\)

OpenStudy (solomonzelman):

Just shifting the linr up

OpenStudy (solomonzelman):

*line

OpenStudy (anonymous):

Alright! Thanks!